OpenStudy (anonymous):

Find the slope of y=-5-5x^2 at the given point (3,-50). Write the equation of the tangent line at the given point.

OpenStudy (jdoe0001):

\(\bf y=-5-5x^2\qquad \cfrac{dy}{dx} = -10x\) if you set x = 3, I'd think that'd be your slope

OpenStudy (anonymous):

My professor gave an example for the class to work with. The example used the equation y=x^2 and gave point (2,4). In the end, its slope was, h+4. Here is the example's process: secant slope= dy/dx = ((2+h)^2-2^2)/h So when I tried to use that process, my slope came out to be h+15. Is that correct?

OpenStudy (anonymous):

Oh I see, what I found was only the secant slope not the slope of the curve. But then what do I do with the secant slope to find the slope of the curve?

hartnn (hartnn):

to get slope of curve from secant slope , we put the limit as h-> 0

hartnn (hartnn):

how did u get h+15 ?

OpenStudy (anonymous):

Sorry, I'm still confused.

OpenStudy (anonymous):

Using the way the example did it, by doing ((3+h)^2-3^2)/h , I got h+15

hartnn (hartnn):

slope of curve \(\Large lim {h \rightarrow 0} \dfrac{f(x+h)-f(x)}{h}\)

hartnn (hartnn):

3 ? isn't the Q have 5 ?

hartnn (hartnn):

-5- 5(3+h)^2 - (-5-3^2) will be the numerator, right ?

OpenStudy (anonymous):

Oh! You're right, I was using the examples function y=x^2 rather than y=−5−5x^2.

hartnn (hartnn):

so, try with -5-x^2 now ?

OpenStudy (anonymous):

would the equation be (((-5-5x^2)+h)-(-5-5x^2))/h ? Because I ended up with 1.

hartnn (hartnn):

we replace 'x' by 'x+h' so, the 1st term would be -5- (x+h)^2 [and not ((-5-5x^2)+h) ] got this ?

OpenStudy (anonymous):

Ok, then wouldn't the first term be -5-5(x^2) ?

hartnn (hartnn):

why would u say so ? its -5 - 5(x+h)^2 and 2nd term = -5-5x^2 subtract those

hartnn (hartnn):

f(3+h)-f(3) = -5- 5(3+h)^2 - (-5-3^2) = ... ?

OpenStudy (anonymous):

These types of questions always get to me. I see now. Could you also just find (f(x+h)−f(x))/h (now I know the right way to do it) and then substitute in the 3 from that result?

OpenStudy (anonymous):

And the 3 was from the x value, right?

hartnn (hartnn):

yes, 3 is from the x-value of the point f(3+h)-f(3) = -5- 5(3+h)^2 - (-5-3^2) = -5 - 5 (h^2+6h+9) +5+9 simplify this simple algebraic expression ?

hartnn (hartnn):

= -5 - 5 (h^2+6h+9) +5+9 = -5h^2 -30h = h (-5h-30) this factored out h, cancels out with the h in the denominator , right ? so you are left with -5h-30 just put h=0 here to get the slope of the curve :)

hartnn (hartnn):

and ask if any more doubts....

OpenStudy (anonymous):

Thank you! :) I've looked all over the place trying to answer this question and it wasn't as hard as I thought.

hartnn (hartnn):

correct! quite simple :) welcome ^_^